OpenStudy (anonymous):

Price of 15 markers box=$12.70. Price of 42 markers box=$31.60 Assuming that there are no taxes what would be the price of a box of 50 markers?

OpenStudy (anonymous):

can anyone please help. Is this a linear relationship problem?

OpenStudy (anonymous):

Call x the price of one marker. Call b the cost of handling Box's price = 12.70 = 15x + b (1) Box's price = 31.60 = 42x + b (2) (2) - (1) -> 18.90 = 27x --> x = 18.90/27 = 0.7 Find b: 12.70 = 15*(0.7) + b -> b = 12.70 - 10.50 = 2.2 The 50 marker box's price: 50*(0.7) + 2.2 = 35 + 2.2 = $37.20

OpenStudy (anonymous):

Thanx a ton! Is just choosing "nest response give u a medal....pls let me know